LOCATING ANISOTROPIES IN ELECTRICAL IMPEDANCE TOMOGRAPHY

ERKKI J. SOMERSALO

Abstract. In this article, we consider the problem of finding the support of an inhomogenous possibly anisotropic inclusion in a background of constant electric conductivity from the electrical impedance tomography data at the boundary of a bounded body. The article discusses the linear sampling method applied to this problem. A practical algorithm for solving this problem is suggested.